OpenStudy (anonymous):

find the exact value of each expression. 1. sin 50° cos 40° + sin 40° cos 50° 2. (tan pi/3 + tan pi/6) / (1- tan pi/3 tan pi/6)

OpenStudy (anonymous):

1. sin (50+40)= sin 90

OpenStudy (anonymous):

@babybaby baby can you do this ?

Parth (parthkohli):

\[\sin(a + b) = \sin a\cos b + \sin b \cos a\]

OpenStudy (anonymous):

@jasonxx i dont really understand the topic

Parth (parthkohli):

Oh, it's Trigonometry.

OpenStudy (anonymous):

it says use the expansion of sin (A+B)= sin A cos B+ cos A sin B where A= 50, B=40 clear baby?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ill try it

OpenStudy (anonymous):

cool :)

hartnn (hartnn):

standard formula : \(\huge tan (A+B)=\frac{tan A+tan B}{1-tanA\: tanB}\) put A = pi/3, B = pi/6 \(\huge tan (\pi/3+\pi/6)=\frac{tan (\pi/3)+tan (\pi/6)}{1-tan(\pi/3)\: tan(\pi/6)}\) so \(\huge\frac{tan (\pi/3)+tan (\pi/6)}{1-tan(\pi/3)\: tan(\pi/6)}=tan(\pi/2)=\infty\) did u get this ?

Directrix (directrix):

Sine of a Sum of Two Angles ------------ sin(α + β) = sinα cos β + cos α sin β. = sin 50° cos 40° + sin 40° cos 50° Decide which is alpha and which is beta, add, and take the sine of the sum.